# Optimization of foreground moment deprojection for semi-blind CMB polarization reconstruction

### v1

@@ 4.2 Optimization of moment deprojection
The improvement induced by combining the various layers of optimization in ocMILC, including the selection of moments and their number, pivot values, and deprojection coefficients, is demonstrated by comparing the power spectrum of ocMILC residuals with those obtained when optimization is exclusively applied to specific layers. This comparison is performed with the following intermediate ocMILC pipelines using the example of the PICO d7 data set (similar results are obtained for other models):
@@
The full ocMILC pipeline, integrating all layers of optimization, has been applied to all PICO $B$-mode data sets, encompassing diverse foreground scenarios outlined in Section 2. The angular power spectra of ocMILC residuals are compared with those of NILC for all foreground scenarios in Figure 6. For the d1, d4, and d7 models, ocMILC yields significantly lower foreground residuals across all angular scales compared to NILC, with only a moderate increase in noise contamination relative to NILC. These improvements in terms of residual foreground contamination, regardless of the foreground model, can be attributed to the several layers of optimization in ocMILC: multiclustering from GNILC, varying number and set of deprojected moments, varying pivot parameters, and varying deprojection coefficients, across the sky and angular scales. For the most challenging foreground model d12, which includes multi-layer dust components along the line-of-sight and spectral decorrelation [29], the reduction of residual foreground contamination is more modest than in other foreground scenarios but still outperforms NILC across angular scales overall. Across all the considered data sets, the reduced residual foreground contamination of the ocMILC CMB $B$-mode solution is achieved with only a moderate noise penalty relative to NILC, thanks to the information provided by the GNILC diagnosis. These results showcase the flexibility of ocMILC in effectively handling deviations of thermal dust emission from a simple MBB through optimized moment deprojection.
To further assess the quality of the CMB $B$-mode reconstruction, we examine the impact of NILC and ocMILC residuals on the estimation of the tensor-to-scalar ratio, $r$. We estimate the tensor-to-scalar ratio from the recovered $B$-mode power spectra using the inverse-Wishart log-likelihood [61]:
Here $\hat{C}_{\ell}=A_{\textrm{lens}}C_{\ell}^{\textrm{lens}}+C_{\ell}^{\textrm{fgds}}+C_{\ell}^{\textrm{noi}}$ represents the angular power spectrum of the output CMB $B$-mode map, averaged over all realizations, with $C_{\ell}^{\textrm{lens}}$ the lensing power spectrum, $A_{\textrm{l}}$ its residual amplitude after some expected level of delensing, $C_{\ell}^{\textrm{fgds}}$ and $C_{\ell}^{\textrm{noi}}$ the foreground and noise residual power, while $C_{\ell}(r)=r\,C_{\ell}^{\,r=1}+A_{\textrm{lens}}C_{\ell}^{\textrm{lens}}+C_{\ell}^{\textrm{noi}}$ represents our theoretical power spectrum model, where $C_{\ell}^{\,r=1}$ is the primordial CMB $B$-mode power spectrum for tensor perturbations with $r=1$. We adopt $A_{\textrm{lens}}=0.27$ to align with the expected $73\%$ delensing capability of PICO [30]. As our simulations do not include primordial tensor perturbations, the best-fit value of $r$ is only driven by residual foreground contamination, while its uncertainty is influenced by residual lensing, noise, and foregrounds. While the likelihood function in Equation 4.13 is not exact for a masked sky as it neglects correlations between different multipoles induced by masking, we ensure that these off-diagonal terms of the covariance matrix are negligible by binning the angular power spectrum with a constant binning scheme of $\Delta\ell=10$.
@@
The resulting $r$-posteriors for NILC and ocMILC across the four distinct foreground scenarios are displayed in the right panels of Figure 6. Biased estimates of the tensor-to-scalar ratio ($r$) are reported when the value $r=0$ falls outside the $68\%$ confidence level. Notably, ocMILC, in contrast to NILC, produces unbiased posteriors for $r$ across all considered foreground scenarios even for such large fractions of sky. This bias removal is characterized by both a shift in the distribution’s peak towards lower values of $r$ due to reduced foreground contamination and an increase in the distribution’s width due to increased noise. The biased $r$-distribution observed in NILC results from more pronounced residual Galactic contamination across all multipoles. While previous PICO forecasts in the literature [23, 30] demonstrated that NILC and cMILC can achieve unbiased recovery of $r$ for a $50\%$ sky fraction with most foreground models, our results highlight that over larger sky fractions ($70\%$) these methods, without optimization, exhibit biases in $r$ due to lingering foreground residuals. In contrast, ocMILC remains unbiased over such extended sky fractions for all foreground models. It is interesting to note that our results, along with those from [30], exhibiting poorer constraints on $r$ for d1 and d12 compared to d4 and d7 foreground scenarios, align with the GNILC diagnosis in Section 3 which indeed revealed lower foreground complexity in the d4 and d7 data sets. For all considered Galactic models, except the intricate d12 model, ocMILC achieves upper bounds on the tensor-to-scalar ratio that align with the sensitivity targeted by PICO with the adopted baseline instrumental configuration ($r<0.0004$ at $95\%$ CL). In Appendix A, we demonstrate on the d1 data set that ocMILC, when applied with the more sensitive PICO CBE instrumental configuration, meets the specified requirement of $r<0.0002$ at $95\%$ CL for this particular configuration.

### v2

@@ 4.2.2 CMB $B$-mode reconstruction
The improvement induced by combining the various layers of optimization in ocMILC is demonstrated by comparing the power spectrum of ocMILC residuals with those obtained when optimization is exclusively applied to specific layers. This comparison is performed using the example of the PICO d7 data set (similar results are obtained for other models) and considering the following intermediate ocMILC pipelines:
@@
The full ocMILC pipeline, integrating all layers of optimization, has been applied to all PICO $B$-mode data sets, encompassing diverse foreground scenarios outlined in Section 2. The angular power spectra of ocMILC residuals are compared with those of NILC for all foreground scenarios in Figure 6. For the d1, d4, and d7 models, ocMILC yields significantly lower foreground residuals across all angular scales compared to NILC, with only a moderate increase in noise contamination relative to NILC. These improvements in terms of residual foreground contamination, regardless of the foreground model, can be attributed to the several layers of optimization in ocMILC: multiclustering from GNILC, varying number and set of deprojected moments, varying pivot parameters, and varying deprojection coefficients, across the sky and angular scales. For the most challenging foreground model d12, which includes multi-layer dust components along the line-of-sight and spectral decorrelation [29], the reduction of residual foreground contamination is more modest than in other foreground scenarios but still outperforms NILC across angular scales overall. For all the considered data sets, the reduced residual foreground contamination of the ocMILC CMB $B$-mode solution is achieved with only a moderate noise penalty relative to NILC, thanks to the information provided by the GNILC diagnosis. These results showcase the flexibility of ocMILC in effectively handling deviations of thermal dust emission from a simple MBB through optimized moment deprojection.
@@
The resulting $r$-posteriors for NILC and ocMILC across the four distinct foreground scenarios are displayed in the right panels of Figure 6. Biased estimates of the tensor-to-scalar ratio ($r$) are reported when the value $r=0$ falls outside the $68\%$ confidence level. Notably, ocMILC, in contrast to NILC, produces unbiased posteriors for $r$ across all considered foreground scenarios even for such large fractions of sky. This bias removal is characterized by both a shift in the distribution’s peak towards lower values of $r$ due to reduced foreground contamination and an increase in the distribution’s width due to increased noise. The biased $r$-distribution observed in NILC results from more pronounced residual Galactic contamination across all multipoles. While previous PICO forecasts in the literature [23, 30] demonstrated that NILC and cMILC can achieve unbiased recovery of $r$ for a $50\%$ sky fraction with most foreground models, our results highlight that over larger sky fractions ($70\%$) these methods, without optimization, exhibit biases in $r$ due to lingering foreground residuals. In contrast, ocMILC remains unbiased over such extended sky fractions for all foreground models. It is interesting to note that our results, along with those from [30], exhibiting poorer constraints on $r$ for d1 and d12 compared to d4 and d7 foreground scenarios, align with the GNILC diagnosis in Section 3.2.1 which indeed revealed lower foreground complexity in the d4 and d7 data sets. For all considered Galactic models, except the intricate d12 model, ocMILC achieves upper bounds on the tensor-to-scalar ratio that align with the sensitivity targeted by PICO with the adopted baseline instrumental configuration ($r<0.0004$ at $95\%$ CL).